OpenStudy (anonymous):

please help is this absolutely or cond convergent sigma (-1)^n-1 arctan(1/n)/n^2

OpenStudy (watchmath):

The term of the absolute series is \(|\arctan(1/n)|/n^2<\frac{\pi/2}{n^2}\). But the series \(\sum 1/n^2\) is convergent (since it is a p-series with p=2), so the absolute series converges. Then the series is absolutely convergent.

OpenStudy (anonymous):

how did you come up with pi/2/x^2

OpenStudy (watchmath):

the range of the arctan function is between -pi/2 and pi/2

OpenStudy (anonymous):

and why did you put over x^2

OpenStudy (watchmath):

well I just you that arctan (1/n) < pi/2. The n^2 is already on the denominator on the first place.

OpenStudy (anonymous):

so would you always compare it to pi/2 divided by denominator given in the problem

OpenStudy (anonymous):

for tan

OpenStudy (watchmath):

I wouldn't say always. I just see the opportunity that if we compare with pi/2 then I can have some conclusion. It is possible for a problem to have an arctan bu we don't compare with the pi/2.

OpenStudy (anonymous):

and if there isnt anything in the denominator of the given problem you would compare it to pi/2

OpenStudy (watchmath):

No, I won't if there is nothing on the bottom, I can compare arctan(1/n) < pi/2. But the series \(\sum pi/2\) is divergent. So comparing with pi/2 doesn't give me any conclusion.

OpenStudy (anonymous):

okay what would you comoare arcsin (1/n) to and 1-cos1/n to

OpenStudy (watchmath):

the term of the series only arcsin(1/n) ?

OpenStudy (anonymous):

yah

OpenStudy (watchmath):

Just to make sure. So your series is \(\sum_{n=1}^\infty \arcsin(1/n)\)? I can't think how to do this right away...

OpenStudy (anonymous):

yah

OpenStudy (watchmath):

ok, arcsin is an increasing function and for positive x we have x > sin x Apply the arcsin arcsin x > x It follows that arcsin(1/n) > 1/n But the harmonic series \(\sum 1/n\) is divergent Hence \(\sum \arcsin(1/n)\) is divergent as well.

OpenStudy (anonymous):

what if it was just sin

OpenStudy (watchmath):

sin what?

OpenStudy (anonymous):

if instead of arc sin it was sin would you still make the same comparison

OpenStudy (watchmath):

you mean sin(1/n) ?

OpenStudy (anonymous):

yah

OpenStudy (watchmath):

well we can't compare to 1/n since 1/n > sin(1/n) So I don't know the answer yet.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

also if you were given sigma 1-cos(1/n) what would you compare it to

OpenStudy (watchmath):

That is actually a nice problem. I'll post it as a new question so everybody can give a response (BTW it is convergent)